6. Team A has probability 2/3 of winning whenever it plays.suppose A plays four Games.find the probability p that A wins more than half of its gamed.


1
Expert's answer
2022-02-24T15:38:52-0500

Let "X=" the number of games won: "X\sim Bin (n, p)."

Given "p=2\/3, n=4, q=1-p=1\/3."


"P(X\ge3)=P(X=3)+P(X=4)"

"=\dbinom{4}{3}(\dfrac{2}{3})^3(\dfrac{1}{3})+\dbinom{4}{4}(\dfrac{2}{3})^4=\dfrac{16}{27}"